OpenStudy (anonymous):

The midpoint of line UV is (5,-11) , the coordinates of one endpoint are U(3,5) , find the coordinates for V ? Is V = (7,-17) ?

OpenStudy (anonymous):

Maaan nobody can answer this question , thanks for nothing geeks

OpenStudy (zaarin):

V=(10/3,-22/5)

OpenStudy (anonymous):

Well thats wrong so thanks for nothing

OpenStudy (zaarin):

X=10-3=7 and Y=-22+5=-17 so. V is (7,-17)